# FedHERO: A Federated Learning Approach for Node Classification Task on Heterophilic Graphs

### v1

@@ 3.4 Computational and Communication Overhead
Let $M$ denote the total number of clients, $E$ represent the number of local training epochs, and $L$ signify the number of layers in the model architecture. For simplicity, let $d$ denote the larger dimension between the node feature dimension $d_{x}$ and the embedding feature dimension $d_{\text{hidden}}$. $n_{\text{link}},\theta^{d},d_{\text{dg}}$, and $d_{\text{rw}}$ regard method-specific parameters. In Table 2, we compute the additional overhead caused by the FGL algorithms compared to FedAvg McMahan et al. (2017).
@@
GCFL Xie et al. (2021) $\mathcal{O}((M(Ld^{2})^{2}+MlogM)/E)$ $\mathcal{O}(1)$
FedLit Xie et al. (2023) $\mathcal{O}(2\lvert\mathcal{E}\rvert n_{link}dI+\lvert\mathcal{V}\rvert d^{2}+2n_{link}^{2}dI/E)$ $\mathcal{O}((n_{link}-1)Ld^{2})$
FedPub Baek et al. (2023) $\mathcal{O}(Ld^{2}+Md/E)$ $\mathcal{O}(1)$
FedStar Tan et al. (2023) $\mathcal{O}(2Ld(d\lvert\mathcal{V}\rvert+\lvert\mathcal{E}\rvert))$ $\mathcal{O}(1)$
FedSage Zhang et al. (2021) $\mathcal{O}(\lvert\mathcal{V}\rvert(\theta^{d}+Ld^{2}))$ $\mathcal{O}(\lvert\mathcal{V}\rvert d+Ld^{2})$
@@ 4.1 Experimental Settings
Datasets. We carry out experiments using well-known heterophily datasets such as Squirrel and Chameleon, derived from Wikipedia datasets Rozemberczki et al. (2021), along with the Actor dataset Tang et al. (2009a) and the Flickr dataset Zeng et al. (2019). Given these datasets’ inherent incompatibility with the FGL framework, we adapt them into a federated context using two prevalent graph partitioning methods: the METIS algorithm Karypis & Kumar (1997) and the Louvain algorithm Blondel et al. (2008). This approach enables us to create semi-synthetic federated graph learning datasets. As detailed in Appendix A, the graph partitioning process preserves the heterophilic properties of the graphs. Additionally, Appendix B demonstrates significant differences in the neighbor distributions of same-class nodes across clients. This highlights the challenges in FGL that we discussed in the Introduction. Following Baek et al. (2023) and Zhang et al. (2021), we assume no overlapping nodes are shared across data owners. For further details on this process and the overlapping scenario, please refer to Appendix A and C, respectively.
We also include three graph datasets: syn-cora Zhu et al. (2020), ieee-fraud Howard et al. (2019), and credit Yeh & Lien (2009). These datasets naturally lend themselves to subdivision into multiple subgraphs (e.g., distinct subgraphs within the ieee-fraud dataset correspond to different product transaction records), and we denote them as real-world datasets. The statistics and description of the datasets are summarized in Appendix A.
Hyper-Parameter Settings. The architecture of the GNN models consists of a feature projection layer, two graph convolutional layers, and a classifier layer. We employ Adam optimizer Kingma & Ba (2014) with a learning rate of 0.005. We apply a consistent hyperparameter setting in our main results. Specifically, we set $\alpha=0.2$, $\mu=\lambda=0.1$, $k=20$, and $N_{H}=4$. In supplementary material, we provide an analysis of hyperparameters, indicating that optimizing these values could lead to further improvements in performance. The training process involves 200 communication rounds, with each local training epoch lasting one iteration. The nodes are randomly distributed into five groups. For each experiment, one group is chosen as the test set, while three of the remaining groups are utilized for training and one for validation. We perform five experiments on each dataset and report the models’ average performance. For the ieee-fraud and credit datasets, we present the mean test Area Under the Curve (AUC), accounting for the imbalanced labels. For the remaining datasets, we use mean test accuracy across clients for evaluation. All methodologies are implemented using PyTorch 1.8 and executed on NVIDIA RTX A6000 GPUs. Our code can be found in the supplementary material.
@@
Baselines. We comprehensively compare FedHERO against five baseline methods in federated graph learning. These baselines encompass strategies that focus on (i) promoting collaboration among clients with similar graph distributions (GCFL Xie et al. (2021) and Fedpub Baek et al. (2023)), (ii) sharing structural insights (FedStar Tan et al. (2023)), (iii) producing missing edges between subgraphs (FedSage Zhang et al. (2021)), and (iv) discerning and modeling message-passing for latent link types (FedLit Xie et al. (2023)). To ensure the completeness of our experiments, we also include the standard FL method FedAvg McMahan et al. (2017).
@@ 4.2 Evaluation on Semi-synthetic Datasets
We also observed that FedAvg often underperforms compared to local training in many experiments. This suggests that heterophilic graphs are more likely to exhibit distinct neighbor distribution patterns across them, which can negatively impact FedAvg’s performance. Notably, FedPub and GCFL demonstrate superior overall performance compared to FedAvg. Both encourage clients to collaborate with those who have similar model gradients or outputs. During optimization, the GNN learns to capture graph information, including neighbor distribution patterns. Therefore, similar model gradients or outputs indicate similar neighbor distribution patterns or tendencies captured by the model. As previously discussed, aggregating models with similar tendencies enhances the performance of the aggregated model. The overall poor performance of FedLit emphasizes that the proposed unique challenge is distinct from the one mentioned in FedLit.
In our investigation of FedHERO’s performance with heterophilic data, we also observed significant improvements in datasets characterized by high heterophily, particularly in the Squirrel and Chameleon datasets. These datasets have homophily ratios of $0.22$ and $0.25$, respectively, as reported in Mao et al. (2023), with lower ratios indicating greater heterophily. In comparison to Flickr (homophily ratio: 0.32 Kim & Oh (2022)), FedHERO’s enhancements are more pronounced. For example, under Louvain partitioning, FedHERO demonstrates improvements of 4.31%, 1.16%, and 0.08% on the Chameleon, Squirrel, and Flickr datasets, respectively, when compared to the best-performing baseline. We believe this phenomenon occurs because, as heterophilic ratio increases, the differences in node neighbor distribution patterns across clients’ local graphs become more pronounced, posing greater challenges to existing FGL methods.
@@ 4.3 Evaluation on Real-world Datasets
Figure 3 provides a comprehensive overview of the evaluation results across real-world datasets in FGL settings. FedHERO consistently outperforms all FGL baselines on ieee-fraud and credit datasets. The similar performance between our method and other FGL baselines on ieee-fraud and credit datasets may be attributed to distinct class features that are easily distinguishable (e.g., a linear regression model can yield satisfactory results) Howard et al. (2019) or the low dimensionality of features Yeh & Lien (2009). The superior performance of FedSage on the syn-cora dataset may be attributed to GraphSage’s utilization of a hybrid design, incorporating elements of both ego-embedding and neighbor embedding. Similar findings are also observed in Zhu et al. (2020).
Notably, FedPub and FedStar face challenges in achieving strong performance on the syn-cora dataset. The syn-cora dataset consists of random graphs where nodes of different classes have the same expected degree, and different graphs share the same degree distribution Abu-El-Haija et al. (2019); Zhu et al. (2020). The suboptimal performance of FedStar also demonstrates that relying solely on structural information from the adjacency matrix, while overlooking distinctions between nodes connected by an edge, is insufficient for handling heterophilic graphs. Thus, structure embeddings in FedStar don’t contain distinguishable information across classes. Sharing the structure knowledge doesn’t benefit training GNNs for node classification. On the other hand, FedPub assumes graphs in clients have community structures, and it leverages the community information. However, the assumption does not hold in syn-cora where random graphs are generated independently.
@@ 4.4 FedHERO with Different Structure Learning Methods
In this section, we explore the adaptability of FedHERO to various structure learning techniques. For example, we substitute the similarity function with an attention-based structure learner, denoted as $\textsc{FedHERO}{att}$, and with Graph Attention Networks (GAT) Veličković et al. (2018), denoted as $\textsc{FedHERO}{gat}$. Another variant, $\textsc{FedHERO}{cos}$, uses cosine distance as the similarity function. We further examine FedHERO with more complicated graph learning methods, including GLCN Jiang et al. (2019), GAug Zhao et al. (2021b). We replace the structure learning model in FedHERO with these methods. The experiments are performed on the Squirrel and Chameleon datasets, and the results are presented in Table 4. From the results, we observe the following:
@@
  • Among basic variants, $\textsc{FedHERO}{cos}$ shows slightly better performance than the other two variants, likely because it avoids additional model parameters, thus reducing the risk of overfitting. Additionally, unlike $\textsc{FedHERO}{gat}$, $\textsc{FedHERO}{cos}$ considers the entire subgraph for edge generation, which mitigates the impact of heterophilic graphs on the GNN aggregation mechanism Li et al. (2022); Yang et al. (2022).
@@ 4.5 FedHERO with Various Sharing Mechanisms
We conduct an analysis of various FedHERO variants involving the sharing of different components of the local model, including all model parameters, task model parameters, and none (denoted as $\textsc{FedHERO}_{all}$, $\textsc{FedHERO}_{gcn}$ and $\textsc{FedHERO}_{none}$, respectively). Additionally, we consider whether to incorporate a structure-task model decoupled scheme. For variants without the Dual-Channel (DC), we remove the GNN in the $f_{local}$ channel and feed the learned adjacency matrix and the original matrix into the same GNN instead.
@@ 4.7 Privacy Preservation in FedHERO
FedHERO only transmits the components in the $f_{global}$ channel to the server. This subset contains fewer parameters, thus reducing the volume of sensitive information transmitted. Furthermore, we evaluate the leakage of structural information from FedHERO through the link inference attack (LIA)  Gong & Liu (2018). LIA aims to infer the underlying graph structure from the learned node representations. The LIA attack accuracies on two datasets are presented in Figure 5 (left). In general, LIA struggles to accurately predict the graph structure due to the inherent heterophilic characteristics of the graphs. We can observe that FedHERO offers better privacy protection compared to the FGL baseline approaches by sharing only a portion of the model, and make LIA even harder to predict.
@@ 5 Related Work
Graph Structure Learning. Graph Neural Networks (GNNs) have emerged as a potent tool for analyzing graph-structured data and have found wide applications in various domains Zhao et al. (2021a); Wu et al. (2019); Kipf & Welling (2016); Zhu et al. (2019); Liu et al. (2020); Fan et al. (2019); Luan et al. (2022); Yuan et al. (2022); Dwivedi et al. (2023); Shlomi et al. (2020). In order to overcome the limitations of GNNs in effectively propagating features within observed structures, recent efforts have aimed at simultaneous learning of graph structures alongside the GNN model Jin et al. (2020); Liu et al. (2022b); Li et al. (2024); Liu et al. (2023b); Wu et al. (2022). At the heart of graph structure learning lies an encoding function responsible for modeling optimal graph structures, often represented by edge weights  Luo et al. (2021); Kreuzer et al. (2021); Sun et al. (2022); Wang et al. (2020); Gasteiger et al. (2019); Yu et al. (2021). These edge weights can be defined through pairwise distances or learnable parameters. For example, GAUGM Zhao et al. (2021b) calculates edge weights by taking the inner product of node embeddings, introducing no additional parameters. Franceschi et al. Franceschi et al. (2019) treat each edge as a Bernoulli random variable and optimize graph structures with the GNN. IDGL Chen et al. (2020), to make full use of information from observed structures for structure learning, employs a multi-head self-attention network to represent pairwise relationships between nodes, while Li et al. Li et al. (2018a) adopt a metric learning approach based on the Radial Basis Function (RBF) kernel for a similar goal. Moreover, DGM Kazi et al. (2022) utilizes the Gumbel-Top-k trick to sample edges from a Gaussian distribution and incorporates reinforcement learning, rewarding edges that contribute to correct classifications and penalizing those leading to misclassifications. Additionally, Zhang et al. Zhang et al. (2019) present a probabilistic framework that views the input graph as a random sample from a collection modeled by a parametric random graph model. However, while these methods have shown promise, they assume that training and testing nodes originate from the same graph, and they primarily consider a single graph. In contrast, Zhao et al. Zhao et al. (2023) address graph structure learning in a cross-graph setting and propose a comprehensive framework for learning a shared structure learner that can generalize to target graphs without necessitating re-training.
@@
Federated Graph Learning. Federated Learning (FL) is a paradigm that facilitates model training across decentralized clients while safeguarding privacy Li et al. (2020); Liu et al. (2024); Gafni et al. (2022); Wang et al. (2023); Ye et al. (2023); Lei et al. (2023). As the field of FL evolves, a new specialized branch has emerged to address the unique challenges of graph data, known as Federated Graph Learning (FGL) Liu et al. (2022a). This branch can be categorized into graph-level and node-level methods.
@@
Graph-level FGL methods consider scenarios where each client possesses multiple graph data, as seen in domains like molecular structures. Like traditional FL, it mainly focuses on the issue of non-independent and identically distribution (non-IID) issues across clients Xie et al. (2021); He et al. (2021); Tan et al. (2023); Zheng et al. (2021); Chen et al. (2021a). On the other hand, node-level FGL assumes that each client holds a single graph, with the objective of addressing node classification tasks. In this variant, each client usually possesses a subgraph of the larger graph, introducing a scenario where there are missing links between subgraphs. To address this issue, Zhang et al. introduced FedSage Zhang et al. (2021), which involves collaboratively learning a missing neighbor generator across clients. However, FedSage could raise privacy concerns. In contrast, Baek et al. Baek et al. (2023) utilize random graphs as inputs to compute similarities between clients’ GNNs, subsequently employing these similarities for weighted averaging on the server side. Additionally, Xie et al. Xie et al. (2023) focus on detecting latent link-types during FGL and differentiating message-passing through various types of links using multiple convolution channels. In this study, we address a specific challenge in node-level FGL, namely the distinction in node neighbor distributions across clients. Our approach involves sharing a structure learning model among clients, enabling the learning of message-passing patterns across the graphs. By leveraging this information, we aim to enhance the predictive performance of the local models.

### v2

@@ 3.4 Computational and Communication Overhead
Let $M$ denote the total number of clients, $E$ represent the number of local training epochs, and $L$ signify the number of layers in the model architecture. For simplicity, let $d$ denote the larger dimension between the node feature dimension $d_{x}$ and the embedding feature dimension $d_{\text{hidden}}$. $n_{\text{link}},\theta^{d},d_{\text{dg}}$, and $d_{\text{rw}}$ regard method-specific parameters. In Table 2, we compute the additional overhead caused by the FGL algorithms compared to FedAvg (McMahan et al., 2017).
@@
GCFL (Xie et al., 2021) $\mathcal{O}((M(Ld^{2})^{2}+MlogM)/E)$ $\mathcal{O}(1)$
FedLit (Xie et al., 2023) $\mathcal{O}(2\lvert\mathcal{E}\rvert n_{link}dI+\lvert\mathcal{V}\rvert d^{2}+2n_{link}^{2}dI/E)$ $\mathcal{O}((n_{link}-1)Ld^{2})$
FedPub (Baek et al., 2023) $\mathcal{O}(Ld^{2}+Md/E)$ $\mathcal{O}(1)$
FedStar (Tan et al., 2023) $\mathcal{O}(2Ld(d\lvert\mathcal{V}\rvert+\lvert\mathcal{E}\rvert))$ $\mathcal{O}(1)$
FedSage (Zhang et al., 2021) $\mathcal{O}(\lvert\mathcal{V}\rvert(\theta^{d}+Ld^{2}))$ $\mathcal{O}(\lvert\mathcal{V}\rvert d+Ld^{2})$
@@ 4.1 Experimental Settings
Datasets. We carry out experiments using well-known heterophily datasets such as Squirrel and Chameleon, derived from Wikipedia datasets (Rozemberczki et al., 2021), along with the Actor dataset (Tang et al., 2009a) and the Flickr dataset (Zeng et al., 2019). Given these datasets’ inherent incompatibility with the FGL framework, we adapt them into a federated context using two prevalent graph partitioning methods: the METIS algorithm (Karypis & Kumar, 1997) and the Louvain algorithm (Blondel et al., 2008). This approach enables us to create semi-synthetic federated graph learning datasets. As detailed in Appendix A, the graph partitioning process preserves the heterophilic properties of the graphs. Additionally, Appendix B demonstrates significant differences in the neighbor distributions of same-class nodes across clients. This highlights the challenges in FGL that we discussed in the Introduction. Following Baek et al. (2023) and Zhang et al. (2021), we assume no overlapping nodes are shared across data owners. For further details on this process and the overlapping scenario, please refer to Appendix A and C, respectively.
We also include three graph datasets: syn-cora (Zhu et al., 2020), ieee-fraud (Howard et al., 2019), and credit (Yeh & Lien, 2009). These datasets naturally lend themselves to subdivision into multiple subgraphs (e.g., distinct subgraphs within the ieee-fraud dataset correspond to different product transaction records), and we denote them as real-world datasets. The description of the datasets is summarized in Appendix A.
Hyper-Parameter Settings. The architecture of the GNN models consists of a feature projection layer, two graph convolutional layers, and a classifier layer. We employ Adam optimizer (Kingma & Ba, 2014) with a learning rate of 0.005. We apply a consistent hyperparameter setting in our main results. Specifically, we set $\alpha=0.2$, $\mu=\lambda=0.1$, $k=20$, and $N_{H}=4$. In the supplementary material, we provide an analysis of hyperparameters, indicating that optimizing these values could lead to further improvements in performance. The training process involves 200 communication rounds, with each local training epoch lasting one iteration. The nodes are randomly distributed into five groups. For each experiment, one group is chosen as the test set, while three of the remaining groups are utilized for training and one for validation. We perform five experiments on each dataset and report the models’ average performance. For the ieee-fraud and credit datasets, we present the mean test Area Under the Curve (AUC), accounting for the imbalanced labels. For the remaining datasets, we use mean test accuracy across clients for evaluation. All methodologies are implemented using PyTorch 1.8 and executed on NVIDIA RTX A6000 GPUs. Our code can be found in the supplementary material.
@@
Baselines. We comprehensively compare FedHERO against five baseline methods in federated graph learning. These baselines encompass strategies that focus on (i) promoting collaboration among clients with similar graph distributions (GCFL (Xie et al., 2021) and Fedpub (Baek et al., 2023)), (ii) sharing structural insights (FedStar (Tan et al., 2023)), (iii) producing missing edges between subgraphs (FedSage (Zhang et al., 2021)), and (iv) discerning and modeling message-passing for latent link types (FedLit (Xie et al., 2023)). To ensure the completeness of our experiments, we also include the standard FL method FedAvg (McMahan et al., 2017).
@@ 4.2 Evaluation on Semi-synthetic Datasets
We also observed that FedAvg often underperforms compared to local training in many experiments. This suggests that heterophilic graphs are more likely to exhibit distinct neighbor distribution patterns across them, which can negatively impact FedAvg’s performance. Notably, FedPub and GCFL demonstrate superior overall performance compared to FedAvg. Both encourage clients to collaborate with those who have similar model gradients or outputs. During optimization, the GNN learns to capture graph information, including neighbor distribution patterns. Therefore, similar model gradients or outputs indicate similar neighbor distribution patterns or tendencies captured by the model. As previously discussed, aggregating models with similar tendencies enhances the performance of the aggregated model.
In our investigation of FedHERO’s performance with heterophilic data, we also observed significant improvements in datasets characterized by high heterophily, particularly in the Squirrel and Chameleon datasets. These datasets have homophily ratios of $0.22$ and $0.25$, respectively, as reported in (Mao et al., 2023), with lower ratios indicating greater heterophily. In comparison to Flickr (homophily ratio: 0.32 (Kim & Oh, 2022)), FedHERO’s enhancements are more pronounced. For example, under Louvain partitioning, FedHERO demonstrates improvements of 4.31%, 1.16%, and 0.08% on the Chameleon, Squirrel, and Flickr datasets, respectively, when compared to the best-performing baseline. We believe this phenomenon occurs because, as heterophilic ratio increases, the differences in node neighbor distribution patterns across clients’ local graphs become more pronounced, posing greater challenges to existing FGL methods.
@@ 4.3 Evaluation on Real-world Datasets
Figure 3 provides a comprehensive overview of the evaluation results across real-world datasets in FGL settings. FedHERO consistently outperforms all FGL baselines on ieee-fraud and credit datasets. The similar performance between our method and other FGL baselines on ieee-fraud and credit datasets may be attributed to distinct class features that are easily distinguishable (e.g., a linear regression model can yield satisfactory results) (Howard et al., 2019) or the low dimensionality of features (Yeh & Lien, 2009). The superior performance of FedSage on the syn-cora dataset may be attributed to GraphSage’s utilization of a hybrid design, incorporating elements of both ego-embedding and neighbor embedding. Similar findings are also observed in (Zhu et al., 2020).
Notably, FedPub and FedStar face challenges in achieving strong performance on the syn-cora dataset. The syn-cora dataset consists of random graphs where nodes of different classes have the same expected degree, and different graphs share the same degree distribution (Abu-El-Haija et al., 2019; Zhu et al., 2020). The suboptimal performance of FedStar also demonstrates that relying solely on structural information from the adjacency matrix, while overlooking distinctions between nodes connected by an edge, is insufficient for handling heterophilic graphs. Thus, structure embeddings in FedStar don’t contain distinguishable information across classes. Sharing the structure knowledge doesn’t benefit training GNNs for node classification. On the other hand, FedPub assumes graphs in clients have community structures, and it leverages the community information. However, the assumption does not hold in syn-cora where random graphs are generated independently.
@@ 4.4 FedHERO with Different Structure Learning Methods
In this section, we explore the adaptability of FedHERO to various structure learning techniques. For example, we substitute the similarity function with an attention-based structure learner, denoted as $\textsc{FedHERO}{att}$, and with Graph Attention Networks (GAT) (Veličković et al., 2018), denoted as $\textsc{FedHERO}{gat}$. Another variant, $\textsc{FedHERO}{cos}$, uses cosine distance as the similarity function. We further examine FedHERO with more complicated graph learning methods, including GLCN (Jiang et al., 2019), GAug (Zhao et al., 2021b). We replace the structure learning model in FedHERO with these methods. The experiments are performed on the Squirrel and Chameleon datasets, and the results are presented in Table 4. From the results, we observe the following:
@@
  • Among basic variants, $\textsc{FedHERO}{cos}$ shows slightly better performance than the other two variants, likely because it avoids additional model parameters, thus reducing the risk of overfitting. Additionally, unlike $\textsc{FedHERO}{gat}$, $\textsc{FedHERO}{cos}$ considers the entire subgraph for edge generation, which mitigates the impact of heterophilic graphs on the GNN aggregation mechanism (Li et al., 2022; Yang et al., 2022).
@@ 4.5 FedHERO with Various Sharing Mechanisms
We conduct an analysis of various FedHERO variants involving the sharing of different components of the local model, including all model parameters, task model parameters, and none (denoted as $\textsc{FedHERO}_{all}$, $\textsc{FedHERO}_{gcn}$, and $\textsc{FedHERO}_{none}$, respectively). Additionally, we consider whether to incorporate a structure-task model decoupled scheme. For variants without the Dual-Channel (DC), we remove the GNN in the $f_{local}$ channel and feed the learned adjacency matrix and the original matrix into the same GNN.
@@ 4.7 Privacy Preservation in FedHERO
FedHERO only transmits the components in the $f_{global}$ channel to the server. This subset contains fewer parameters, thus reducing the volume of sensitive information transmitted. Furthermore, we evaluate the leakage of structural information from FedHERO through the link inference attack (LIA)  (Gong & Liu, 2018). LIA aims to infer the underlying graph structure from the learned node representations. The LIA attack accuracies on two datasets are presented in Figure 5 (left). In general, LIA struggles to accurately predict the graph structure due to the inherent heterophilic characteristics of the graphs. We can observe that FedHERO offers better privacy protection compared to the FGL baseline approaches by sharing only a portion of the model, and make LIA even harder to predict.
@@ 5 Related Work
Graph Structure Learning. Graph Neural Networks (GNNs) have emerged as a potent tool for analyzing graph-structured data and have found wide applications in various domains (Zhao et al., 2021a; Wu et al., 2019; Kipf & Welling, 2016; Zhu et al., 2019; Liu et al., 2020; Fan et al., 2019; Luan et al., 2022; Yuan et al., 2022; Dwivedi et al., 2023; Shlomi et al., 2020). In order to overcome the limitations of GNNs in effectively propagating features within observed structures, recent efforts have aimed at simultaneous learning of graph structures alongside the GNN model (Jin et al., 2020; Liu et al., 2022b; Li et al., 2024; Liu et al., 2023b; Wu et al., 2022). At the heart of graph structure learning lies an encoding function responsible for modeling optimal graph structures, often represented by edge weights  (Luo et al., 2021; Kreuzer et al., 2021; Sun et al., 2022; Wang et al., 2020; Gasteiger et al., 2019; Yu et al., 2021). These edge weights can be defined through pairwise distances or learnable parameters. For example, GAUGM (Zhao et al., 2021b) calculates edge weights by taking the inner product of node embeddings, introducing no additional parameters.  Franceschi et al. (2019) treat each edge as a Bernoulli random variable and optimize graph structures with the GNN. IDGL (Chen et al., 2020), to make full use of information from observed structures for structure learning, employs a multi-head self-attention network to represent pairwise relationships between nodes, while Li et al. (Li et al., 2018a) adopt a metric learning approach based on the Radial Basis Function (RBF) kernel for a similar goal. Moreover, DGM (Kazi et al., 2022) utilizes the Gumbel-Top-k trick to sample edges from a Gaussian distribution and incorporates reinforcement learning, rewarding edges that contribute to correct classifications and penalizing those leading to misclassifications. Additionally, Zhang et al. (2019) present a probabilistic framework that views the input graph as a random sample from a collection modeled by a parametric random graph model. However, while these methods have shown promise, they assume that training and testing nodes originate from the same graph, and they primarily consider a single graph. In contrast, Zhao et al. (2023) address graph structure learning in a cross-graph setting and propose a comprehensive framework for learning a shared structure learner that can generalize to target graphs without necessitating re-training.
@@
Federated Graph Learning. Federated Learning (FL) is a paradigm that facilitates model training across decentralized clients while safeguarding privacy (Li et al., 2020; Liu et al., 2024; Gafni et al., 2022; Wang et al., 2023; Ye et al., 2023; Lei et al., 2023). As the field of FL evolves, a new specialized branch has emerged to address the unique challenges of graph data, known as Federated Graph Learning (FGL) (Liu et al., 2022a). This branch can be categorized into graph-level and node-level methods.
@@
Graph-level FGL methods consider scenarios where each client possesses multiple graph data, as seen in domains like molecular structures. Like traditional FL, it mainly focuses on the issue of non-independent and identically distribution (non-IID) issues across clients (Xie et al., 2021; He et al., 2021; Tan et al., 2023; Zheng et al., 2021; Chen et al., 2021a). On the other hand, node-level FGL assumes that each client holds a single graph, with the objective of addressing node classification tasks. In this variant, each client usually possesses a subgraph of the larger graph, introducing a scenario where there are missing links between subgraphs. To address this issue, Zhang et al. introduced FedSage (Zhang et al., 2021), which involves collaboratively learning a missing neighbor generator across clients. However, FedSage could raise privacy concerns. In contrast, Baek et al. (2023) utilize random graphs as inputs to compute similarities between clients’ GNNs, subsequently employing these similarities for weighted averaging on the server side. Additionally, Xie et al. (2023) focus on detecting latent link-types during FGL and differentiating message-passing through various types of links using multiple convolution channels. In this study, we address a specific challenge in node-level FGL, namely the distinction in node neighbor distributions across clients. Our approach involves sharing a structure learning model among clients, enabling the learning of message-passing patterns across the graphs. By leveraging this information, we aim to enhance the predictive performance of the local models.